OpenStudy (anonymous):

Motion Problems - D = R x T S = Skateboard B = Bicycle Together, a skateboard and a bicycle cost $199.00. The skateboard costs $51.00 less than a bicycle. How much does each cost?

OpenStudy (anonymous):

well... total cost = 199 skateboard costs (x-51), where x = cost of bicycle. x + (x-51) = 199 this statement says "cost of bicycle + cost of skateboard = the total cost of them together" just use some algebra to solve for x 2x - 51 = 199 2x = 250 x = 125 now figure out using x what the individual cost is skateboard cost: x-51 125-51 skateboard costs $74 obviously bicycle is just x, so it costs $125 125+74 = 199.

OpenStudy (anonymous):

IDK where d*rt is supposed to be used or why you originally posted it... but i dun see how that is relevant to this type of problem eh (this isn't really a "motion" problem, there is no movement).

OpenStudy (anonymous):

just to keep you up to date with the formula DRT

hero (hero):

I came here expecting to help someone with a "motion" problem. Instead, I get a problem related to the cost of skateboards and bikes. Guess I'll have to keep looking :P